OpenStudy (anonymous):

if csc (t) = 3 and t is in the second quadrant, evaluate sin (pi - t). i know sin (t) = 1/3, and that the arcsin (1/3) gives me t, so i can just plug it in to sin (pi - t) to find that it is 1/3. positive since it is sin in quad 2. but there must be a way to do it by hand (not using arcsin function from calculator).

OpenStudy (anonymous):

doesn't matter what quadrant you are in it is always true that \[\sin(\pi-x)=\sin(x)\]

OpenStudy (anonymous):

because pi is 180?

OpenStudy (anonymous):



OpenStudy (anonymous):

you should put degrees out of your mind when working with radians. like going to another country where they use different money. don't convert, just work with what you have

OpenStudy (anonymous):

sorry the picture is so childish. the point is that you are the same height on the unit circle, so the y - coordinate is the same

OpenStudy (anonymous):

hmm.. i guess that makes sense. but could i get a step by step explanation of the drawing. its kind of hard to understand just showing it all at once :(